OpenStudy (anonymous):

Compare the following functions: f(x) = −4 sin(3x − π) + 2 g(x) graph of a quadratic with points at 1, 0 and 3, negative 4 and 5, 0 h(x) x y −2 14 −1 9 0 6 1 5 2 6 3 9 4 14 Which function has the smallest minimum? f(x) g(x) h(x) All three functions have the same minimum @Nnesha @Directrix @zepdrix @iambatman @whpalmer4 @Loser66 @Kainui

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Helloo!!

OpenStudy (misty1212):

the maximum of sine is \(-\) so the minimum of \(-4\sin(whatever)+2\) is \(-4+2=-2\)

OpenStudy (anonymous):

right

OpenStudy (misty1212):

the minimum value of h looks like it is 5

OpenStudy (anonymous):

yes it is 5

OpenStudy (misty1212):

the quadratic function g looks like it has a minimum of -5

OpenStudy (misty1212):

oh no it say -4 not -5 you wrote it in words right?

OpenStudy (misty1212):

minus 4 that is the smallest minimum

OpenStudy (anonymous):

so the answer will be g(x)?

OpenStudy (misty1212):

yes i would say \(g(x)\) has the smallest minimum of -4

OpenStudy (anonymous):

Thank you soo muchh!! @misty1212

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

do you mind helping me on another one? <3 @misty1212

OpenStudy (misty1212):

lol no of course not

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

here is the question

OpenStudy (anonymous):

What is the arc length when Θ = 3 pi over 5 and the radius is 7 cm? 21 pi over 5cm 12 pi over 5cm 6 pi over 5cm 3 pi over 35cm @misty1212

OpenStudy (misty1212):

\[\frac{3\pi}{5}\times 7\]

OpenStudy (misty1212):

i guess \(\frac{21\pi}{5}\)

OpenStudy (misty1212):

you ok now ? \(\color\magenta\heartsuit\)

OpenStudy (anonymous):

yes thx!!

OpenStudy (anonymous):

that @misty1212 seems pretty sharp

OpenStudy (anonymous):

what do u mean?

OpenStudy (anonymous):

gave you some good looking answers is all how is it going today?

OpenStudy (anonymous):

good thank you!! how about yu?

OpenStudy (anonymous):

sad today

OpenStudy (anonymous):

my cat ran away, i hope she comes back

zepdrix (zepdrix):

@satellite73 OH NO!! Mrs kitty whiskers ran away? That's so sad. :c Cat's. are. awesome.

OpenStudy (anonymous):

awwwwwwwwww oh nooooo!!! hopefully she does! @satellite73

OpenStudy (anonymous):

@zepdrix can you plz help me?!

zepdrix (zepdrix):

with what? 0_o what's going on?

OpenStudy (anonymous):

just a math question

zepdrix (zepdrix):

what question 0_o

OpenStudy (anonymous):

What is the sine value of 2 pi over 3? negative 1 over 2 1 over 2 negative square root 3 over 2 square root 3 over 2

zepdrix (zepdrix):

2pi/3 is located somewhere over here.

zepdrix (zepdrix):

The sine of that angle corresponds to the y coordinate.

zepdrix (zepdrix):

Do you have a chart or something you can look at? This mostly just has to do with memorization :d

OpenStudy (anonymous):

no.. sorry :/

OpenStudy (anonymous):

OMG!! what the heck is THAT?!

zepdrix (zepdrix):

you haven't seen that yet? 0_o mmm weird...

zepdrix (zepdrix):

See the dot for the angle 2pi/3? The y-coordinate corresponds to your sine value.

OpenStudy (anonymous):

ok so the answer is -1/2?

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

that's the x coordinate :O

OpenStudy (anonymous):

oh okay then it must be sqrt 3/2

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (anonymous):

thanks!!